OpenStudy (anonymous):

5(4X-3)-5(3x-4)=4x+8 X=?

OpenStudy (anonymous):

20x-15-15x+20=4x+8 5x+5=4x+8 x=3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You are welcome! ;)